# Optimal time estimation and the clock uncertainty relation for stochastic processes

### v1

@@ IV.2 Accuracy
Increased accuracy therefore comes at the cost of reduced resolution, such that at best the accuracy scales with the inverse resolution, $\mathcal{N}\sim\nu^{-1}$. This is a tighter special case of the general trade-off found in Ref. [7], which states that any (classical or quantum) clock’s accuracy scales at most quadratically with the inverse resolution. That is, $\mathcal{N}\leq\mathcal{A}^{2}_{\rm max}/\nu^{2}$, where $\mathcal{A}_{\rm max}$ is the mean rate at which monitored transitions can occur maximised over all states of the clock, e.g. in our setup $\mathcal{A}_{\rm max}=\max_{\mathbf{p}}\sum_{(\sigma\to\mu)\in\aleph}R_{\mu\sigma}p_{\sigma}$.
@@ IV.3 Allan variance
Using the techniques of Ref. [44], an explicit expression for the Allan variance can be found in terms of the current autocorrelation function $F(\tau)={\rm Cov}[I(t+\tau),I(t)]$. Full details are given in Appendix J, while here we simply note the short- and long-time limits
@@
where $D$ is the diffusion coefficient and $K=\sum_{\mu,\sigma}w_{\mu\sigma}^{2}J_{\mu\sigma}$ is a weighted dynamical activity, which both depend on the choice of weights $\vec{w}$. For long times, minimising the Allan variance is equivalent to minimising $D$ and therefore the BLUE is optimal, yielding the long-time scaling $\sigma_{y}^{2}(T)\sim\mathcal{T}/T$. Any other solution to Eq. (28) would similarly minimise the long-time Allan variance. For short times, however, the Allan variance is minimised by choosing weights that extremise $K$ instead. Taking into account the unbiased constraint, $\vec{w}\cdot\vec{J}=1$, this extremum is achieved by the uniform weighting $w_{\mu\sigma}=\mathcal{A}^{-1}$, i.e. $\Theta=N/\mathcal{A}$ where $N$ is the total number of jumps and $\mathcal{A}$ is the dynamical activity (see Appendix J). This yields the minimal Allan variance $\sigma_{y}^{2}(T)\sim 1/(\mathcal{A}T)$ at short times. Notice that the same estimator saturates the CRB at short times, as discussed in Sec. II.4.
@@
To illustrate the crossover from this short- to the long-time regime, we plot results for a unicyclic clock model in Fig. 8. We allow both clockwise and anticlockwise jumps with disordered rates to ensure that differences between $\mathcal{A}$ and $\mathcal{T}^{-1}$ are visible, but to ensure reproducibility we generate these rates from a deterministic sequence; see the caption of Fig. 8 for details. We see that the uniformly weighted estimator has the smallest Allan variance at short times (red squares in Fig. 8). However, after a timescale comparable to a single jump, $T\sim\mathcal{T}$, this uniform estimator transitions to a suboptimal scaling that is worse than that of the BLUE. Notably, for the BLUE one can show that $\sigma_{y}^{2}(T)=\mathcal{T}/T$ holds exactly at all times (blue circles in Fig. 8). This can be understood due to the delta-correlated nature of the current (cf. Eq. (30)), so that correlations across different averaging periods vanish (see Appendix J)
@@ V Discussion
Stochastic uncertainty relations are often discussed using terminology borrowed from metrology (e.g. precision, SNR, etc.) and they have been derived in many cases using the mathematics of parameter estimation, with the Cramér-Rao bound (CRB) playing a prominent role [72, 73, 71, 29, 36, 74]. Information theory applied to Markov jump processes has also been used to bound the precision of measurements under resource constraints [83, 68, 69, 84]. The clock uncertainty relation (CUR) presented here weaves yet another thread in the tapestry of connections between metrology and stochastic thermodynamics. The CUR states that steady-state fluctuations are bounded by the system’s capacity to measure time precisely. In simple terms, the more frequently a system transitions between its underlying states, the more precisely it can function as a clock, and the optimal time estimators coincide with the observables that fluctuate least. We emphasise that the CUR is a tight bound because it quantifies freneticity via the inverse mean residual time. This contrasts with the kinetic uncertainty relation, which cannot generally be saturated except for in the short-time limit, where it coincides with the CRB for time estimation.
@@
While we remain hopeful that no unfortunate readers will need to use our results on stochastic time estimation to survive a shipwreck, some remarks on their practical implementation are still called for. In particular, it is not necessary to already know the waiting times $\Gamma_{\sigma}^{-1}$ precisely in order to construct an asymptotically optimal time estimator. Such an estimator can be determined experimentally by observing the long-time current statistics to infer $\mathbbm{D}$ and $\vec{J}$ from Eqs. (III.1). An optimal counting observable can then be constructed using any pseudoinverse of $\mathbbm{D}$ to solve Eq. (28). Moreover, while application of Eqs. (III.1) would seem to require a priori knowledge of $t$, this is not the case: one can use a given process, say $N_{12}$, as a “time standard” by setting $J_{12}=1$ (tantamount to a choice of units), and then define all other quantities in terms of ratios $N_{\mu\sigma}/N_{12}$. It is intriguing that any estimator thus obtained is generally not a thermodynamic current in the sense of being odd under time reversal. Indeed, the optimal estimator does not depend on the order of the sequence $\bm{\sigma}$ at all, so it would assign the same estimates to a trajectory and its time reverse. Counter-intuitively, therefore, the most precise estimators of time cannot distinguish the direction of time’s arrow.
@@ Acknowledgements.
[…]
@@
MTM and GTL conceived of the project and developed the mathematical framework. NN and RS established the degeneracy of the diffusion tensor and conjectured the form of the optimal time estimator and the clock uncertainty relation. PPP proved the conjecture using the saddle-point approximation and KP performed analytical and numerical calculations for the two-state model to verify the results. FM developed the concept of resolution and introduced and solved the Erlang clock example. MTM drafted the manuscript and produced the plots with input from all authors, and NN created the illustrations in Fig. 1. All authors contributed substantially to the discussion and interpretation of the results and ideas underpinning this work.
@@ Appendix A Path probabilities and waiting-time distributions
Appendix B Current statistics
B.1 Elementary stochastic currents
@@
where the contribution from ${\rm d}t^{-1}\to\delta(\tau)$ tends to a Dirac delta function in the continuum limit. For $\tau>0$, we have, again using Eqs. (B.1),
@@
B.2 Current statistics of a general counting observable
@@
$D=\lim_{t\to\infty}\frac{{\rm d}}{{\rm d}t}{\rm Var}[\Theta(t)]=\sum_{\alpha,\beta,\mu,\sigma}w_{\alpha\beta}D_{\alpha\beta\mu\sigma}w_{\mu\sigma}\equiv\vec{w}\cdot\mathbbm{D}\cdot\vec{w},$ (80)
where $\Theta(t)=\int_{0}^{t}{\rm d}t^{\prime}\,I(t^{\prime})$ is the corresponding integrated current. We represent “vectors” with $d^{2}$ elements using the notation $\vec{w}$, $\vec{I}(t)$, etc., while the tensors $\mathbbm{D}$ and $\mathbbm{F}$ act as $d^{2}\times d^{2}$ matrices, mapping one such vector onto another.
@@
B.3 Hyperaccurate current and the BLUE

### v2

@@ IV.2 Accuracy
Increased accuracy therefore comes at the cost of reduced resolution, such that at best the accuracy scales with the inverse resolution, $\mathcal{N}\sim\nu^{-1}$. This is a tighter special case of the general trade-off found in Ref. [7], which states that any (classical or quantum) clock’s accuracy scales at most quadratically with the inverse resolution. That is, $\mathcal{N}\leq\mathcal{A}^{2}_{\rm max}/\nu^{2}$, where $\mathcal{A}_{\rm max}$ is the mean rate at which monitored transitions can occur maximised over all states of the clock, e.g. in our setup $\mathcal{A}_{\rm max}=\max_{\mathbf{p}}\sum_{(\sigma\to\mu)\in\aleph}R_{\mu\sigma}p_{\sigma}$, where the maximum is performed over all probability distributions $\mathbf{p}$.
@@ IV.3 Allan variance
Using the techniques of Ref. [46], an explicit expression for the Allan variance can be found in terms of the current autocorrelation function $F(\tau)={\rm Cov}[I(t+\tau),I(t)]$. Full details are given in Appendix L, while here we simply note the short- and long-time limits
@@
where $D$ is the diffusion coefficient and $K=\sum_{\mu,\sigma}w_{\mu\sigma}^{2}J_{\mu\sigma}$ is a weighted dynamical activity, which both depend on the choice of weights $\vec{w}$. For long times, minimising the Allan variance is equivalent to minimising $D$ and therefore the BLUE is optimal, yielding the long-time scaling $\sigma_{y}^{2}(T)\sim\mathcal{T}/T$. Any other solution to Eq. (28) would similarly minimise the long-time Allan variance. For short times, however, the Allan variance is minimised by choosing weights that extremise $K$ instead. Taking into account the unbiased constraint, $\vec{w}\cdot\vec{J}=1$, this extremum is achieved by the uniform weighting $w_{\mu\sigma}=\mathcal{A}^{-1}$, i.e. $\Theta=N/\mathcal{A}$ where $N$ is the total number of jumps and $\mathcal{A}$ is the dynamical activity (see Appendix L). This yields the minimal Allan variance $\sigma_{y}^{2}(T)\sim 1/(\mathcal{A}T)$ at short times. Notice that the same estimator saturates the CRB at short times, as discussed in Sec. II.4.
@@
To illustrate the crossover from this short- to the long-time regime, we plot results for a unicyclic clock model in Fig. 9. We allow both clockwise and anticlockwise jumps with disordered rates to ensure that differences between $\mathcal{A}$ and $\mathcal{T}^{-1}$ are visible, but to ensure reproducibility we generate these rates from a deterministic sequence; see the caption of Fig. 9 for details. We see that the uniformly weighted estimator has the smallest Allan variance at short times (red squares in Fig. 9). However, after a timescale comparable to a single jump, $T\sim\mathcal{T}$, this uniform estimator transitions to a suboptimal scaling that is worse than that of the BLUE. Notably, for the BLUE one can show that $\sigma_{y}^{2}(T)=\mathcal{T}/T$ holds exactly at all times (blue circles in Fig. 9). This can be understood due to the delta-correlated nature of the current (cf. Eq. (30)), so that correlations across different averaging periods vanish (see Appendix L)
@@ V Discussion
Stochastic uncertainty relations are often discussed using terminology borrowed from metrology (e.g. precision, SNR, etc.) and they have been derived in many cases using the mathematics of parameter estimation, with the Cramér-Rao bound (CRB) playing a prominent role [81, 82, 80, 29, 36, 83]. Information theory applied to Markov jump processes has also been used to bound the precision of measurements under resource constraints [86, 77, 78, 98]. The clock uncertainty relation (CUR) presented here weaves yet another thread in the tapestry of connections between metrology and stochastic thermodynamics. The CUR states that steady-state fluctuations are bounded by the system’s capacity to measure time precisely. In simple terms, the less time before a system makes its first jump, the more precisely it can function as a clock at long times, and the optimal time estimators coincide with the observables that fluctuate least. We emphasise that the CUR is a tight bound because it quantifies freneticity via the inverse mean residual time. Crucially, this feature can be exploited to detect the presence of hidden states using the observable jump and waiting-time statistics. This contrasts with the kinetic uncertainty relation, which cannot generally be saturated except for in the short-time limit, where it coincides with the CRB for time estimation.
@@
While we remain hopeful that no unfortunate readers will need to use our results on stochastic time estimation to survive a shipwreck, some remarks on their practical implementation are still called for. In particular, it is not necessary to already know the waiting times $\Gamma_{\sigma}^{-1}$ precisely in order to construct an asymptotically optimal time estimator. Such an estimator can be determined experimentally by observing the long-time current statistics to infer $\mathbb{D}$ and $\vec{J}$ from Eqs. (III.1). An optimal counting observable can then be constructed using any pseudoinverse of $\mathbb{D}$ to solve Eq. (28). Moreover, while application of Eqs. (III.1) would seem to require a priori knowledge of $t$, this is not the case: one can use a given process, say $N_{12}$, as a “time standard” by setting $J_{12}=1$ (tantamount to a choice of units), and then define all other quantities in terms of ratios $N_{\mu\sigma}/N_{12}$. It is intriguing that any estimator thus obtained is generally not a thermodynamic current in the sense of being odd under time reversal. Indeed, the optimal estimator does not depend on the order of the sequence $\bm{\sigma}$ at all, so it would assign the same estimates to a trajectory and its time reverse. Counter-intuitively, therefore, the most precise estimators of time cannot distinguish the direction of time’s arrow.
@@
Note added
Soon after the completion of our work, Macieszczak [111] reported similar results obtained using large-deviation theory. This complementary method allowed CUR-like bounds (referred to in Ref. [111] as “ultimate KUR”) to be derived for counting observables and first-passage times, not only for classical Markov dynamics but also semi-Markov processes, a class which includes many examples of quantum reset clocks studied in the recent literature [1, 4, 5, 6, 7, 112, 113].
@@ Acknowledgements.
We thank Michael J. Kewming for helpful discussions on the Allan variance of counting estimators, and Juan P. Garrahan for insightful correspondence on the manuscript. This research was supported in part by the National Science Foundation under Grants No. NSF PHY-1748958 and PHY-2309135. We specifically acknowledge the KITP program New Directions in Quantum Metrology and the conference Time in Quantum Theory (TiQT) 2023, where some of this work was performed. This project is co-funded by the European Union (Quantum Flagship project ASPECTS, Grant Agreement No. 101080167). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union, Research Executive Agency or UKRI. Neither the European Union nor UKRI can be held responsible for them. KP and PPP acknowledge funding from the Swiss National Science Foundation (Eccellenza Professorial Fellowship PCEFP2_194268). RS acknowledges funding from the Swiss National Science Foundation via an Ambizione grant PZ00P2_185986. MTM is supported by a Royal Society University Research Fellowship.
@@
MTM and GTL conceived of the project and developed the mathematical framework. NN and RS established the degeneracy of the diffusion tensor and conjectured the form of the optimal time estimator and the clock uncertainty relation. PPP proved the conjecture using the saddle-point approximation, and KP performed analytical and numerical calculations for the two-state model to verify the results and numerical simulations to detect a hidden Markov model. FM developed the concept of resolution and introduced and solved the Erlang clock example. MTM drafted the manuscript and produced the plots with input from all authors, and NN created the illustrations in Figs. 1 and 2. All authors contributed substantially to the discussion and interpretation of the results and ideas underpinning this work.
@@ Appendix A Path probabilities and waiting-time distributions
Here, $\overleftarrow{\rm T}$ denotes the time-ordering symbol, which reorders all subsequent time-dependent operators such that their time arguments increase from right to left, i.e.
$\overleftarrow{\rm T}\mathsf{A}(t_{1})\mathsf{B}(t_{2})=\begin{cases}\mathsf{A}(t_{1})\mathsf{B}(t_{2}),&(t_{1}\geq t_{2})\\ \mathsf{B}(t_{2})\mathsf{A}(t_{1}),&(t_{1}<t_{2})\end{cases}.$ (59)
@@
Appendix B Jump steady state and jump-conditioned transition probabilities
In this appendix we provide more details on the jump steady state discussed in Sec. II.2, and the conditional transition probabilities used to define resolution in Sec. IV.1. Suppose that a jump is observed in a given small time interval ${\rm d}t\to 0$. Under steady-state conditions, the probability that the observed jump was $\sigma\to\mu$ is given by
$P(\sigma\to\mu|\,{\rm jump})=\frac{R_{\mu\sigma}p_{\sigma}^{\rm ss}}{\mathcal{A}},$ (73)
where $\mathcal{A}=\sum_{\mu,\sigma}R_{\mu\sigma}p_{\sigma}^{\rm ss}$ is the dynamical activity. This follows from the master equation, which defines the probability to observe a jump $\sigma\to\mu$ in the time interval ${\rm d}t$ as $P(\sigma\to\mu,\,{\rm jump})=R_{\mu\sigma}p^{\rm ss}_{\sigma}{\rm d}t$, such that the total probability for any jump to be observed is $P({\rm jump})=\mathcal{A}{\rm d}t$. Putting these together gives $P(\sigma\to\mu|\,{\rm jump})=P(\sigma\to\mu,\,{\rm jump})/P({\rm jump}).$
Marginalising Eq. (73) over the final state after the jump, $\mu$, then yields
where the jump steady state, $p^{\rm jss}_{\sigma}$, was introduced in Eq. (11). Marginalising Eq. (73) over the initial state $\sigma$ yields the same distribution by stationarity, i.e. $\sum_{\sigma}R_{\mu\sigma}p_{\sigma}^{\rm ss}/\mathcal{A}=p_{\mu}^{\rm jss}$. Therefore, $p_{\mu}^{\rm jss}$ is both the fraction of jumps into and out of state $\mu$. Equivalently, $p_{\mu}^{\rm jss}$ is the frequency of the state $\mu$ within a long sequence $\bm{\sigma}$. The jump steady state is also the stationary distribution of the transition probabilities $\pi(\mu|\sigma)=R_{\mu\sigma}/\Gamma_{\sigma}$, as defined below Eq. (4) in the main text. This means that
$\mathsf{L}=\begin{pmatrix}-2&1&4\\ 1&-4&4\\ 1&3&-8\end{pmatrix},$ (76)
Now suppose that we monitor only those jumps $\sigma\to\mu$ within an alphabet $\aleph$, as discussed above Eq. (50), which defines the resolution. Given that some jump within $\aleph$ is observed, the probability that it was the specific jump $\sigma\to\mu$ is
$P(\sigma\to\mu|\aleph)=\frac{R_{\mu\sigma}p_{\sigma}^{\rm ss}}{\mathcal{A}_{\aleph}},$ (77)
by a straightforward adaptation of Eq. (73). Above, the partial dynamical activity for the alphabet $\aleph$ is defined as $\mathcal{A}_{\aleph}=\sum_{(\sigma\to\mu)\in\aleph}R_{\mu\sigma}p_{\sigma}^{\rm ss}$. For a given generator $\mathsf{L}$ and its associated steady state $p_{\sigma}^{\rm ss}$, $\mathcal{A}_{\aleph}$ depends only on the alphabet $\aleph$. Moreover, since it is a sum of positive terms, decreasing the size of the alphabet can only decrease $\mathcal{A}_{\aleph}$. Therefore, for any other alphabet $\aleph^{\prime}\subseteq\aleph$, we have $\mathcal{A}_{\aleph^{\prime}}\leq\mathcal{A}_{\aleph}$.
Appendix C Distribution of residual time
$W(\tau)=\sum_{\sigma}p_{\sigma}^{\rm jss}W(\tau|\sigma),$ (78)
given some initial distribution $p_{\sigma_{1}}$ and after marginalising over the initial and final states of both jumps. In Eq. (79) we are conditioning on precisely $N=2$ jumps having occurred but the total time is unconstrained. This is known as the $N$-ensemble, as opposed to the $t$-ensemble used in the majority of this work. (See Sec. VI.B of Ref. [46] for a detailed discussion of the $t$- and $N$-ensembles; see also Refs. [115, 116] where they are referred to as the $s$- and $x$-ensembles, respectively.) Stationarity means that the marginal distributions of $\tau_{1}$ and $\tau_{2}$ are identical. This is the case only if we set $p_{\sigma_{1}}=p_{\sigma_{1}}^{\rm jss}$ equal to the jump steady state, whence
This result is valid for any discrete, stationary process where intervals between events are characterised by a distribution $W(\tau)$ with mean $\mathcal{A}^{-1}$ [65, 90]. In particular, it holds for arbitrary non-Markovian dynamics so long as the process is stationary. For the specific case of Markovian jump processes considered here, we find simply
$W_{0}(\tau)=\sum_{\sigma}p_{\sigma}^{\rm ss}W(\tau|\sigma),$ (84)
as expected from first principles.
Appendix D Limiting behaviour of the Fisher information
D.1 Saddle-point approximation
$\sum_{\mu}\frac{n_{\mu}}{\Gamma_{\mu}-z_{\rm sp}}=t,$ (89)
$\partial_{t}\ln P(\bm{\sigma}|t)=-z_{\rm sp}+\mathcal{O}(t^{-1}).$ (91)
Together with Eq. (91), this recovers Eq. (15) in the main text.
D.2 Short-time Fisher information
The score is therefore given to leading order by
Combining Eqs. (93) and (94), we have
$P(1|t)=\mathcal{A}t+\mathcal{O}(t^{2})=1-P(0|t),$ (96)
${\rm E}[N]=\mathcal{A}t+\mathcal{O}(t^{2})={\rm Var}[N].$ (97)
Appendix E Statistics of the asymptotically efficient estimator
$\Theta(\mathbf{n})=\sum_{\mu}\frac{n_{\mu}}{\Gamma_{\mu}},$ (98)
as well as the final expression for the Fisher information in Eq. (17). The derivations in this appendix exploit several results from Appendix G: specifically, the statistics of the elementary counting observables $N_{\mu\sigma}(\bm{\sigma})$, i.e. the number of jumps $\sigma\to\mu$ in a given sequence $\bm{\sigma}$. The reader may thus find it helpful to consult Appendix G before proceeding.
Our starting point is the representation
We also have
where $P(\sigma_{1}\cdots\sigma_{N},\mathbf{t}|\sigma_{0},t)=P(\bm{\sigma},\mathbf{t}|t)/p_{\sigma_{0}}^{\rm ss}$ is the probability for a trajectory, conditioned on the system starting in the state $\sigma_{0}$. Using Eqs. (G.1) and (122), we therefore have
and, following some simple manipulations, the variance
${\rm Var}[\Theta]$ $=\mathcal{T}t+{\rm Var}[\Gamma_{\sigma}^{-1}]$
Here, we define
These expressions are exact for all $t$, but simple approximations are possible in the limit of times that are very short, $\Gamma_{\sigma}t\ll 1$, or very long, $\Gamma_{\sigma}t\gg 1$. For short times, the integral in Eq. (E) is negligible, while for long times we can extend the integration limit to infinity and use Eq. (131). We thus obtain
It is interesting to compare these results with the best unbiased linear estimator (BLUE): a minimal-variance counting observable that is defined in Sec. III.1. As discussed there, and shown in detail in Appendix G.3, the variance of the BLUE is exactly equal to $\mathcal{T}t$ for all times. For short times, therefore, the uncertainty in Eq. (109) is increased relative to the BLUE because $\Theta(\mathbf{n})=\Gamma_{\sigma_{0}}^{-1}\neq 0$ even though no jump has yet occurred. For large times, conversely, the uncertainty is decreased by the same amount since all states along the trajectory are accounted for. Note that this decrease is of order $\mathcal{O}(t^{0})$ and is therefore negligible under the same approximations that lead to Eq. (17). Hence, there is no inconsistency with the CRB within the saddle-point approximation, which neglects subleading corrections in time.
Appendix F Exact solution for the two-state system
To illustrate the estimator $\Theta(\mathbf{n})$ given in Eq. (16) and the validity of the saddle-point approximation in obtaining Eqs. (15) and (17), which was presented in Appendix D.1, we investigate a two-state system. Let $\Gamma_{\mu}$ be the escape rate from the state $\mu=1,2$, such that the transition rates are given by $R_{21}=\Gamma_{1}$ and $R_{12}=\Gamma_{2}$. The steady state of the system reads $p_{1}^{\rm ss}=\Gamma_{2}/(\Gamma_{1}+\Gamma_{2})$, $p_{2}^{\rm ss}=\Gamma_{1}/(\Gamma_{1}+\Gamma_{2})$. For the two-state system, each trajectory $\bm{\sigma}$ is unambiguously determined by the number of visits of each state, $\mathbf{n}=(n_{1},n_{2})$, together with the initial state $\sigma_{0}$.
Starting from Eqs. (71) and (A), we may write
where
In the reminder of this section, we illustrate the validity of the saddle-point approximation used to analytically obtain Eqs. (15) and (17) in the asymptotic limit, which is shown in Appendix D.1. In the asymptotic limit, we have $n_{\mu}\approx n_{\mu}+1$, and thus we set $n_{1}=n_{2}=n$. The sum of residues in Eq. (F) may be expressed as
resulting in
The last term asymptotically vanishes, leading to the expression
In Appendix D.1, the saddle-point approximation was employed to show that in the asymptotic regime $\partial_{t}\ln P(\bm{\sigma}|t)=-z_{\rm sp}$, where the saddle point, $z_{\rm sp}$, is determined by Eq. (89), which may equivalently be expressed as a quadratic equation:
Appendix G Current statistics
G.1 Elementary stochastic currents
@@
where the contribution from ${\rm d}t^{-1}\to\delta(\tau)$ tends to a Dirac delta function in the continuum limit. For $\tau>0$, we have, again using Eqs. (G.1),
@@
G.2 Current statistics of a general counting observable
@@
$D=\lim_{t\to\infty}\frac{{\rm d}}{{\rm d}t}{\rm Var}[\Theta(t)]=\sum_{\alpha,\beta,\mu,\sigma}w_{\alpha\beta}D_{\alpha\beta\mu\sigma}w_{\mu\sigma}\equiv\vec{w}\cdot\mathbb{D}\cdot\vec{w},$ (135)
where $\Theta(t)=\int_{0}^{t}{\rm d}t^{\prime}\,I(t^{\prime})$ is the corresponding integrated current. We represent “vectors” with $d^{2}$ elements using the notation $\vec{w}$, $\vec{I}(t)$, etc., while the tensors $\mathbb{D}$ and $\mathbbm{F}$ act as $d^{2}\times d^{2}$ matrices, mapping one such vector onto another.
@@
G.3 Hyperaccurate current and the BLUE